OpenStudy (anonymous):

What is the property of 3/2+d=d+3/2 0=c+(-c)

OpenStudy (skullpatrol):

The first is the addition property of equality $$a+c=b+c$$ and the second is the property of opposites.$$a+(-a)=0$$